OpenStudy (anonymous):

need to help with an algebra 1 question. posting it below

OpenStudy (anonymous):

2. Use graphing technology to graph the equation for each motor. Set rounding to 5 places. Save a copy of each graph and paste them here or attach them to the assignment. If you are using GeoGebra, follow these instructions to find the vertex and x-intercepts. (15 points total, 5 points each graph) 3. Look at each motor's graph and find the vertex coordinates. Then report them below. (12 points total, 4 points each) Motor A Vertex coordinates: Motor B Vertex coordinates: Motor C Vertex coordinates

OpenStudy (anonymous):

the equations that i need to use are = -16x^2 + 190x + 0 =690 -16x^2 + 210x + 0 =690 -16x^2 + 230x + 0 =690

OpenStudy (anonymous):

im confused on how to do this

OpenStudy (anonymous):

@Hero or anyone pls

OpenStudy (anonymous):

Are you the instagram girl?

OpenStudy (anonymous):

and could u help me on this question

OpenStudy (anonymous):

Is your instagram name awk0perf?

OpenStudy (anonymous):

Oh, that's cool. Good luck with that.

OpenStudy (anonymous):

@dmezzullo

OpenStudy (dmezzullo):

Are posting exam questions?

OpenStudy (anonymous):

no its an assignment not an exam. if u could could u walk me through the process @dmezzullo

OpenStudy (anonymous):

that would be great

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

If you posted it in a more understandable way, I could maybe help.